if 60 answers are wrong out of 250 questions, what is the percentage of right answers?
Answer by Maths68(1474)

Wrong answers = 60
Right answers = 250-60 = 190
Percentage of right answers = (190/250)*100
Percentage of right answers = (19/25)*100
Percentage of right answers = (0.76)*100
Percentage of right answers = 76%